OpenStudy (anonymous):

please answer with solution if the measure of an angle of a triangle is 10degrees less than the 1st angle and the 3rd angle is 10degrees less than twice the first, then the triangle is classified as what kind of triangle according to its angles?

OpenStudy (zzr0ck3r):

50 40 90 this help?

OpenStudy (zzr0ck3r):

draw a triangle with those angles and you will get it

OpenStudy (anonymous):

do you have solution about this?

OpenStudy (zzr0ck3r):

what do we know about a triange with a angle that is 90 dagress?

OpenStudy (anonymous):

Rectangle.

OpenStudy (zzr0ck3r):

we are told its A triangle

OpenStudy (zzr0ck3r):

really we want x + (x-10) + (2x-10) = 180

OpenStudy (lgbasallote):

let x be the first angle the second angle is 10 degrees less than the first so the second angle is x - 10 the third angle is 10 degrees less than twice the first so the third degree is 2x - 10 since this is a triangle...the total of the measure of the angles is 180 so x + (x-10) + (2x-10) = 180 x + x - 10 + 2x - 10 = 180 combine the terms 4x - 20 = 180 add 20 to both sides 4x = 200 divide both sides by 4 so x = 50<--this is the first angle the second angle is 10 degrees less so it's 40 then the third angle is 10 degrees less than twice of 50 so the third angle is 90 so the angles are 50, 40 and 90 i decided to post the solution since no one bothered to give it..

OpenStudy (zzr0ck3r):

x+x-10+2x-10 = 180 4x-20 = 180 4x = 200 x = 50 so we know angle one is 50 and we know angle two is angle 1 -10 thus 40 and we know angle three is 2*angle 1 minus 10 thus 90 a right triangle

OpenStudy (anonymous):

thanks guys...